OpenStudy (anonymous):

HELPPP How can you solve the equation 4^-x = 2^x + 3 graphically?

Directrix (directrix):

Draw the graphs of y = 4 - x and y = 2x +3 on the same set of axes and see where they intersect. @maria_brianna

Directrix (directrix):

Do you need to draw the graphs or is there more to this problem?

OpenStudy (anonymous):

yes it asks to make tables to find the solution to 4^-x =2^x+3 and to take the integer values of x between -3 and 3 @Directrix

OpenStudy (phi):

is it \[ 4^{-x} = 2^{x+3} \]??

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

@phi

OpenStudy (phi):

they want you to find the "y" values for the equation \[ y = 4^{-x} \] and the equation \[y= 2^{x+3} \]

OpenStudy (phi):

for x= -3, -2, -1, 0 , 1 , 3 I got that from ***make tables to find the solution to 4^-x =2^x+3 and to take the integer values of x between -3 and 3***

OpenStudy (phi):

for the first equation \[ y = 4^{-x}\] do you know how to figure out y when x is -3 ? you replace the x with -3: \[ y= 4^{-(-3)} \] minus a minus is a plus, so that simplifies to \[ y= 4^3\] can you figure out 4^3 (= 4*4*4, or use ^ on your calculator) what do you get ?

OpenStudy (anonymous):

i got 64

OpenStudy (phi):

now make a table with two columns. the left column is labeled x, the right col is y in the x column, put in -3, and next to it in the y column put 64

OpenStudy (phi):

now try to figure out y when x is -2 can you do that ?

OpenStudy (anonymous):

i got 0.0625

OpenStudy (phi):

can you show the steps?

OpenStudy (anonymous):

i just typed 4^-2 in my calculater

OpenStudy (phi):

oh. that is close. but the equation is 4^(-x)

OpenStudy (anonymous):

oh my bad its 16

OpenStudy (phi):

ok. put -2, 16 in your table now do x=-1 which means do 4^(-(-1)) which is 4^1 (which you can do in your head, right ?)

OpenStudy (anonymous):

its 4

OpenStudy (phi):

now x=0 you can do this in your head, because anything to the zero power is ??

OpenStudy (anonymous):

1

OpenStudy (phi):

now do x=1, 2, 3 we will now start getting fractions (or decimals)

OpenStudy (anonymous):

i got the same numbers tho

OpenStudy (phi):

are you forgetting the equation uses -x (not just x)?

OpenStudy (anonymous):

oh lol yea

OpenStudy (anonymous):

for 1 got 0.25, for 2 i got 0.0625, for 3 i got 0.015625

OpenStudy (phi):

yes, or 1/4, 1/16 , 1/64 now you should have the table they want for the first equation. now we do the same thing for the other equation \[ y = 2^{x+3} \]

OpenStudy (phi):

what y do you get for x= -3 ?

OpenStudy (anonymous):

i got 64

OpenStudy (phi):

start by replace x with -3 in the equation

OpenStudy (phi):

so you get \[ y= 2^{-3+3} \] where I live, -3+3 is 0

OpenStudy (phi):

and 2^0 is ??

OpenStudy (anonymous):

1

OpenStudy (phi):

so in this new table (label it the y= 2^(x+3) table) you get x=-3, y =0 ok ?

OpenStudy (phi):

now find y when x= -2

OpenStudy (anonymous):

so theres gonna be 2 tables?

OpenStudy (phi):

yes

OpenStudy (anonymous):

so for starters of the table it would be the same numbers ?? -3 between 3?

OpenStudy (phi):

the same x numbers (the left column) the right column will have different y numbers (because it's a different equation) the idea is that somewhere (at some x) the y values are the same for both equations at that x. that is the spot we are trying to find.

OpenStudy (phi):

if the curves never cross we won't find a solution. but they are being nice... they are more or less telling us to expect the curves to cross somewheres between x=-3 and x=3

OpenStudy (anonymous):

oh okk

OpenStudy (phi):

now find y when x= -2

OpenStudy (anonymous):

i got 3.25

OpenStudy (phi):

can you show your steps?

OpenStudy (anonymous):

i plugged 2^-2+3 in the calculator

OpenStudy (phi):

order of operations means the calculator did exponents first it did \[ 2^{-2} \] and then added 3 but you want to do \[ 2^{-2+3} \] which is a different thing. first do the -2+3 you can put parens in for the calculator. or first figure out -2+3, and use that number for the exponent.

OpenStudy (anonymous):

so its just 1?

OpenStudy (phi):

for the exponent y= 2^(-2+3) = 2^1 = ??

OpenStudy (anonymous):

oh 2

OpenStudy (phi):

so add -2, 2 to your table now do x= -1

OpenStudy (anonymous):

wait what was the one for x collom -3 ?

OpenStudy (phi):

y= 2^(-3+3) = 2^(?)

OpenStudy (anonymous):

1

OpenStudy (phi):

for x=-3 you get 2^0 = 1

OpenStudy (anonymous):

and -1=4

OpenStudy (phi):

I would say -1, 4 (the numbers are not equal, but -1 and 4 "go together" )

OpenStudy (anonymous):

ok

OpenStudy (phi):

and for x=-2 you found y= 2 -2, 2 now do x= 0 (getting back on track, I hope)

OpenStudy (anonymous):

its 8?

OpenStudy (phi):

yes 2^(0+3) = 2^3 = 2*2*2 = 8 so 0, 8 are the entries in your table for x, y now x=1

OpenStudy (anonymous):

x=1 is 16 x=2 is 32 x=3 is 64

OpenStudy (phi):

yes. now we look at the two tables. at x=-3 the first equation's y is "above" (bigger) than the y value of the 2nd equation.

OpenStudy (phi):

at x=3 the first equation's y is "below" (less than) the y value of the 2nd equation. this means the equations "crossed" somewhere. any idea what x it was?

OpenStudy (anonymous):

no

OpenStudy (phi):

look at your two tables. do you see the y's getting smaller in the first table? do you see the y's getting bigger in the 2nd table (as we move from x=-3 to x=+3)? where are the y's the same?

OpenStudy (anonymous):

ohhhhhh theyre the same at -1

OpenStudy (phi):

which means \[ 4^{-x} = 2^{x+3} \] is true at x= -1 we solved the problem!

OpenStudy (anonymous):

yaayy thank you so much!!

OpenStudy (phi):

most times, we will find the point where the first equation's y is bigger than the 2nd's and at the next x , the y value is smaller. that means the "answer" is an x in between the two we used. We would have to work harder, looking for the answer in between those two x's but here, it was easy (thank goodness) I hope some of this was helpful. Good luck!

OpenStudy (anonymous):

it was thanks!!